Find the frequency of revolution of the electron in the classical model of the hydrogen atom. In what region of the spectrum are electromagnetic waves of this frequency?

1
Expert's answer
2021-11-30T17:45:10-0500



"\nu=\frac{v_n}{2\pi r_n}"


"mv_nr_n=\frac{nh}{2\pi}\to v_n=\frac{nh}{2\pi mr_n}"


"m\frac{v_n^2}{r_n}=\frac{q^2}{4\pi\epsilon_0r_n^2}\to r_n=\frac{h^2\epsilon_0}{\pi m q^2}\cdot n^2"


"v_n=\frac{nh}{2\pi mr_n}=\frac{nh}{2\pi m}\frac{\pi m q^2}{h^2\epsilon_0n^2}=\frac{q^2}{2nh\epsilon_0}"


"\nu_n=\frac{v_n}{2\pi r_n}=\frac{q^2}{2\pi \cdot2nh\epsilon_0}\cdot \frac{\pi m q^2}{h^2\epsilon_0n^2}="


"=\frac{q^4 m}{4h^3\epsilon_0^2}\cdot\frac{1}{n^3}"


For "n=1" "\nu_1=\frac{(1.6\cdot10^{-19})^4\cdot9.1\cdot10^{-31}}{4\cdot (6.62\cdot10^{-34})^3\cdot(8.85\cdot10^{-12})^2}\cdot\frac{1}{1^3}\approx6.56\cdot10^{15}\ (Hz)"


which corresponds to ultraviolet region.